# Analyzing the Effect of Linguistic Similarity on Cross-Lingual Transfer: Tasks and Experimental Setups Matter

### v1

@@ 1 Introduction
For many of the world’s languages, the available data to train machine learning models is scarce. If data is available, it is often only enough for a validation set, raising the question of how to select the training set. Two approaches are intuitive: (i) based on linguistic similarity features: find the training data in a language that is linguistically as close as possible to the target language, where “linguistically close” is defined by one or multiple linguistic similarity measures, or (ii) based on dataset dependent features: find a dataset in another language that has useful features to the target language, such as a high word or $n$-gram overlap. Naturally, these two types of approaches are not mutually exclusive.
@@
However, it is unclear which of these features are most important to select the source language for cross-lingual transfer, despite previous work focusing on this question (see §2): earlier studies often contradict each other, and therefore leave a number of important avenues for improvement. For example, prior work often lacks a large representation of languages and language families, as well as NLP tasks, and sometimes relies on synthetically constructed datasets. Furthermore, popular out-of-the-box measures for linguistic similarity are sometimes intransparent or faulty (Toossi et al., 2024; Khan et al., 2024; §3.3), and correlations between different similarity measures are often not taken into account. Summarizing, how findings from prior work generalize across a larger variety of languages and tasks, and how we should interpret different similarity scores, are open questions.
@@
    Different similarity measures are important for different tasks and input representations. For instance, syntactic similarity is most predictive for POS tagging and parsing, whereas trigram overlap is the most important predictor for $n$-gram-based topic classification (§4.3);
@@
    A feature’s predictive performance on a task in one language transfers relatively well to another similar task (§4.4);
@@
    When no information on cross-lingual transfer performance is available for a language for one task, choosing a transfer language based on another similar task is a relatively safe choice (§4.4).
@@ 2 Related Work
Philippy et al. (2023) survey contributing factors for linguistic similarity. An important take-away from their work is that prior research presents contradicting findings. For example, some studies find lexical overlap to correlate more strongly with token rather than sentence classification tasks Srinivasan et al. (2021), whereas other work finds the opposite Ahuja et al. (2022). Similar contradictory examples are presented for different linguistic similarity metrics, pre-training dataset size, and model architecture. Based on this insight, Philippy et al. make a number of recommendations for follow-up work: (i) focus on real natural languages (as opposed to synthetic ones), (ii) examine the interaction between different contributing factors, (iii) focus on a wide range of languages, (iv) focus on linguistic features when selecting training languages, and (v) focus on generative tasks, given the recent success of generative models. We take these recommendations, specifically focusing on the first four. In our work, we choose to focus on three classification tasks (POS tagging, dependency parsing, and topic classification), due to the availability of train and test data in a large number of languages for these tasks.
We next highlight some of the work on cross-lingual classification tasks that is particularly relevant for our study. de Vries et al. (2022) focus entirely on cross-lingual transfer for POS tagging. They make use of the Universal Dependencies (UD) dataset de Marneffe et al. (2021), and include $65$ languages for training, and $105$ languages for evaluation. Samardžić et al. (2022) also use UD, and focus on dependency parsing with $47$ source and $62$ target languages. Lin et al. (2019) and Xia et al. (2020) also focus on entity linking and machine translation besides POS tagging and dependency parsing. Lauscher et al. (2020) include five NLP tasks, with one source language and 8–14 target languages, finding different transfer trends for token- vs. sentence-level tasks. Adelani et al. (2024) include cross-lingual topic classification experiments with four source and around $197$ target languages. We extend their work by including all languages in the dataset in our analysis for topic classification.
An important difference that distinguishes our work from prior work is that we include a substantially larger number of languages in our analysis. Moreover, in comparison to the studies that come closest to our work in terms of number of included languages de Vries et al. (2022); Samardžić et al. (2022) we include a larger variety of tasks, allowing us to draw conclusions across task boundaries.
@@ 3 Methodology
We run transfer experiments on two word-level grammatical tasks with word-level annotations (§3.1) and a sentence-level topic classification task (§3.2). Of the 266 languages in our experiments, 55 training and 84 test languages are shared between all tasks. Appendix §A contains details on all languages and datasets in this study. §3.3 introduces our similarity metrics.
@@ 3.1 Grammatical tasks
We use the UDPipe 2 models (Straka, 2018, 2023) that were trained on UD 2.12 data. Each model is trained on a single treebank. The models have so far only been systematically evaluated on their within-treebank performance, but not cross-lingually.
UDPipe 2 combines multiple types of input representations: monolingual character and word embeddings (trained on the available training treebanks for a given language) and embeddings derived from mBERT Devlin et al. (2019). The models are trained jointly for POS tagging, dependency parsing, lemmatization and morphological feature prediction. We only evaluate on the first two, as many treebanks do not have gold standard labels for the others. UDPipe 2 post-processes the predicted dependencies to ensure that each sentence includes a root dependency that all other nodes (in)directly depend on.
We evaluate POS tagging using accuracy, and dependency parsing using the labeled (LAS) and unlabeled attachment scores (UAS). For LAS, we follow the UD evaluation script [^1] https://github.com/UniversalDependencies/tools/blob/352505c/eval.py and ignore dependency label subtypes.
@@ 3.3 Similarity measures
We include a range of dataset-dependent and -independent similarity measures, which are similar to measures used in related work (§2). This includes more intuitive measures like syntactic similarity or word overlap, and some measures for which we did not assume high correlations with the output (phonological, phonetic and geographic similarity). We include those for comparability with other work and as baselines for the other similarity comparisons.
@@
Grambank Skirgård et al. (2023a, b) encodes grammatical information for several thousand languages. Its 195 grammatical features are chosen to allow almost no logical dependencies between the values of different features, i.e. the value of one feature does not logically entail the value of another.
@@
We use lang2vec’s location information (each language is represented as a vector of distances to a number of points on the Earth’s surface) and calculate the Euclidean distances between language vectors. We do not use lang2vec’s precomputed distances, as the package erroneously returns the maximum distance for some languages, regardless of the comparison. We define geographic proximity (geo) as 1 minus the distance.
@@
We count the number of sentences in each training dataset (size). For the topic classification task, this is trivial as each language has the same number of training samples (701 sentences).
@@
4 Results and Analysis
We first compare the performance within and across languages for the different tasks (§4.1). Next, we consider the effect of the writing system (§4.2). We then investigate which factors correlate with overall transfer performance (§4.3). Finally, we explore what this means for selecting a source language for cross-lingual transfer (§4.4).
4.1 Performance within and across languages
The different tasks and models show different transfer patterns, both regarding how well transfer works overall and which language pairs form clusters.
4.1.1 Comparison within tasks
@@
Unsurprisingly, the within-language performances are much higher than the cross-lingual performances. [^4] We focus on the language level for most of our analyses. For UD, which has nearly twice as many datasets as languages, we find that trends are very stable across datasets of the same language, whereas writing systems play an important role for the SIB-200 data. More details are in §E.2. Table 7 in §E shows the mean scores within and across languages and datasets for all tasks, which we summarize below.
Grammatical tasks
Take-away: Both within-language and cross-lingual performance are higher for POS tagging than for parsing. For POS tagging, the within-language performance is high (mean accuracy: $93.4\%$, compared to 43.0% cross-lingually). Scores are lower for parsing (mean within-language LAS: 78.4% (UAS: 84.8%) vs. mean cross-lingual LAS: 20.6% (UAS: 36.8%)).
@@
Our POS tagging results are similar but not identical to the ones that de Vries et al. (2022) obtained on UD v2.8 with finetuned XLM-R models for 65 training and 105 test languages. When we select the v2.8 languages and use language-level averages, our results are similar (Pearson’s $r$ and Spearman’s $\rho=0.73$, $p<0.0001$) – while the model choice (incl. input representations) matters for the resulting patterns, many transfer trends are similar.
Topic classification
Take-away: The monolingual $n$-gram-based models have higher within-language but lower cross-lingual performance than the multilingual model.
The mean accuracy of topics-base is $66.7$% for within-language evaluation and 20.1% cross-lingually. Topics-translit achieves the same mean within-language accuracy, but performs slightly better cross-lingually (22.5%). [^5] Its within-dataset performance is slightly lower than for topics-base (69.4% vs. 70.3%), likely due to some language-specific information being lost when diacritics are removed. This would be compensated in the within-language performance by improved transfer between datasets of languages that originally had different scripts.
@@
The model using mBERT representations has the lowest within-language accuracy (60.1%). We hypothesize that this due to high [UNK] token rates for some languages that were not part of mBERT’s pretraining data. [^6] Bagheri Nezhad and Agrawal (2024) similarly find that the performance of multilingual models that are monolingually trained and evaluated on SIB-200 relates to how well a language is represented in a model’s pretraining data. The model’s mean cross-lingual performance, however, benefits from the multilingual input embeddings: It is about twice as high as for the monolingual models (42.6%). Generally, transfer works well if both training and test languages are in mBERT’s pretraining data (or closely related to a language that is). The average accuracy is 68.3% if both languages were included in mBERT’s pretraining data, 46.5% if only the target language was included, 36.1% if only the source language was included, and 33.2% if neither was included. This is consistent with results by Adelani et al. (2024), who finetuned four XLM-Rlarge models on one high-resource language each and found their transfer results to be very similar to each other.
@@
Figure 2: Different experiments produce different transfer patterns. NLP transfer results for all combinations of training (columns) and test languages (rows). The three heatmaps for the grammatical tasks are sorted in the same order, and the three heatmaps with the topic classification results are sorted in the same order. The darker diagonal shows the within-language scores. Large, labelled heatmaps are in Appendix §E.
4.1.2 Comparison across tasks
@@
In Figure 2, we plot heatmaps of the transfer results for different tasks. To save space, we omit the language labels (which can be found in §E.1).
Take-away: The transfer patterns are similar across related tasks, but we also observe some different patterns. We compute correlation scores between different tasks, which we summarize below (details in §F.1).
@@
The results of the grammatical tasks are highly correlated with each other, the correlations between other task results are weaker. Because of the very high correlation between the parsing measures LAS and UAS (Pearson’s $r=0.95$), we focus on LAS in the remainder of the paper as the more extensive of the two: LAS not only measures whether the dependency arcs are correct, but also whether they were labelled correctly.
Even when using exactly the same underlying data and models, evaluating on a related task results in transfer trends that are similar, but not entirely the same: The parsing (LAS) and POS tagging results produced by identical UDPipe models (except for the classification heads) are correlated with $r=0.86$.
@@
The preprocessing of the input data (e.g., transliteration) and especially the choice of input representation (mono- vs. multilingual) affects the transfer trends: The results of the two $n$-gram-based topic classifiers are highly correlated with each other ($r=0.68$), but not with the results of the mBERT-based set-up ($r=0.28$ and $0.36$).
Set-ups that involve multilingual representations are more highly correlated. For the languages that appear both in UD and SIB-200, the results of the grammatical tasks are most strongly correlated with those of the mBERT-based topic classification models ($r=0.64$ for POS tagging, $r=0.58$ for LAS).
4.2 Effect of writing system
Take-away: Transfer between datasets with the same writing systems generally works better than between different scripts, however this is in part due to the language selection rather than the scripts themselves. Two-sample Kolmogorov–Smirnov tests indicate that the results within vs. across scripts come from different distributions (p-values all <0.0001; statistics: 0.51 for topics-base, 0.26 for topics-mbert, 0.16 for POS accuracy, 0.18 for LAS). However, the results of topics-translit for datasets that were in the same vs. different scripts before transliteration also come from different distributions (p <0.0001, statistic: 0.37), indicating that at least for SIB-200, the combinations of languages usually associated with the scripts alone already make an important difference.
Nonetheless, writing systems still play a role. Eight of the languages in SIB-200 come in two script versions each. For all topic classification models, the within-language scores are higher when the writing systems are identical than when they are different (details in §E.3).
4.3 Correlations between similarity measures and transfer results
@@
We next calculate the correlations between the similarity measures and the NLP task results. To account for differences between test sets, we calculate correlations (Pearson’s r) between similarity measures and transfer results on a test language level. We compare the performance of different NLP systems on the same test language, but we do not compare the performance of a single system across multiple test sets. This decision is motivated especially by the challenges of comparing parsing performance across treebanks. Sentence length influences parsing difficulty as it determines the available search space (cf. McDonald and Nivre, 2011, Choi et al., 2015), and morphological differences between languages hamper the comparability of parsers Nivre and Fang (2017).
@@
We analyze overall correlation scores by averaging across test languages. We treat correlation coefficients with p-values of at least 0.05 as 0, and exclude items where we could not calculate similarity scores due to missing linguistic information. We use language-level averages so that those languages with multiple training and test datasets do not have artificially high correlations, and languages with multiple test datasets do not have greater influence on the overall averages. However, if we instead consider treebank-level correlations, we observe that test sets that belong to the same language show very similar correlation patterns and scores.
@@
The correlations between task results and similarity measures vary across our experiments. Figure 3 shows the correlations, which we summarize below. Unaggregated correlations and 95% confidence intervals of the mean scores can be found in §F.3.
@@
Figure 3: Mean correlation scores between task results and similarity measures. “Word*” refers to overlap between words for the UD tasks (wor), trigrams (tri) for topics-base and topics-translit, and subword tokens (swt) for topics-mbert. Dotted/dashed lines are added for intelligibility.
Common trends
@@
Training dataset size does not matter much in our experiment: the training size is the same across all topic classification experiments, and the correlations between training set size and performance in the grammatical tasks are close to zero. The phonological and phonetic features (pho, inv) also generally have low correlation scores. [^7] Philippy et al. (2023) find that phonological similarity is more important for token- than sentence-level tasks. This is different in our study, where the correlations with inv are higher for the $n$-gram-based topic classification models. We assume that this is due to their being moderately correlated with string similarity measures like trigram overlap, which is in turn strongly correlated with the topic classification results.
Parsing
@@
Take-away: The strongest predictor for parsing performance is syntactic similarity (syn) as determined by lang2vec ($r$avg_syn$=0.57$), followed by the similarity of Grambank features (gb) ($r$avg_gb$=0.42$). These correlations are even stronger when we only consider the test languages that mBERT was pretrained on ($r$avg_syn$=0.69$ and $r$avg_gb$=0.60$, respectively; Appendix §F.3).
POS tagging
Take-away: The POS tagging outputs show correlation patterns similar to the ones for parsing, albeit weaker. Although the correlation strengths are generally weaker for POS tagging, e.g., the strongest predictor is lang2vec’s syntactic similarity (syn) with $r$avg_syn$=0.37$), word overlap (wor) is as good a predictor for POS accuracy as for LAS ($r$avg_wor=0.33 and $r$avg_wor=0.34).
Again, the correlation strengths are higher when only considering test languages that were in mBERT’s pretraining data ($r$avg_syn$=0.46$). The correlation strengths are also higher when we analyze de Vries et al.’s results ($r$avg_syn$=0.46$). We hypothesize that this is due to the smaller amount of test languages in their experiments, a higher proportion of which is in XLM-R’s pretraining data.
Topic classification
@@
Take-away 1: The results of the $n$-gram-based models are most highly correlated with measures of string similarity and lexical similarity. This is expected based on their input representations. For topics-base, the trigram overlap (tri) between training and test data shows the highest correlation ($r$avg_tri=0.65), followed by character overlap (chr) ($r$avg_chr=0.49) and lexical similarity (lex) ($r$avg_lex=0.49). Trigram overlap and lexical similarity (and the highly correlated genetic relatedness (gen)) are also the strongest predictors for topics-translit ($r$avg_tri=0.55, $r$avg_lex=0.61, $r$avg_gen=0.55). However, for this model, character overlap only plays a very small role ($r$avg_chr=0.11), as all datasets use the same character inventory due to the transliteration.
@@
Take-away 2: For the model using mBERT representations (topics-mbert), none of the correlations are strong. The correlations peak at $r$avg_chr=0.27 for character overlap. Model performance depends instead on the inclusion of the source and especially target languages in mBERT’s pretraining data (§4.1.1).
Independent effects
@@
Take-away: The transfer results cannot be predicted by any one factor alone. We fit a linear mixed effects model for each experiment, with the NLP score as the dependent variable and the source and target languages as random effects. The fixed effects are the similarity scores, whether the training and test data use the same writing systems, and whether the test language was included in mBERT’s pretraining data. We use R R Core Team (2024) and the lme4 package Bates et al. (2015). This analysis is carried out on a reduced set of languages, as for many language pairs in our data at least one similarity metric is not defined and we do not perform any data imputation. This analysis shows similar trends to the correlations described above. Additionally, we find that for each experiment, most of the variables considered in our analyses are significant predictors of the NLP score (details in §F.2). This even sometimes applies to measures that capture similarity at the same linguistic level: e.g., both grammatical similarity measures (syn, gb) are independently among the highest predictors for parsing performance.
4.4 Practical implications: Choosing transfer languages based on simple heuristics

### v2

@@ 1 Introduction
For many of the world’s languages, the available data to train machine learning models is scarce. If data is available, it is often only enough for a validation set, raising the question of how to select the training set. Two approaches are intuitive: (i) based on linguistic similarity measures: find the training data in a language that is linguistically as close as possible to the target language, where “linguistically close” is defined by one or multiple linguistic similarity measures, or (ii) based on dataset dependent measures: find a dataset in another language that is similar to the target dataset, e.g., in terms of high word or $n$-gram overlap. Naturally, these two types of approaches are not mutually exclusive.
@@
However, it is unclear which of these measures are most important to select the source language for cross-lingual transfer, despite previous work focusing on this question (see §2): earlier studies often contradict each other, and therefore leave a number of important avenues for improvement. For example, prior work often lacks a large representation of languages and language families, as well as NLP tasks, and sometimes relies on synthetically constructed datasets. Furthermore, popular out-of-the-box measures for linguistic similarity are sometimes intransparent or faulty (Toossi et al., 2024; Khan et al., 2025; §3.3), and correlations between different similarity measures are often not taken into account. Summarizing, how findings from prior work generalize across a larger variety of languages and tasks, and how we should interpret different similarity scores, are open questions.
@@
    For different tasks and input representations, different similarity measures are most predictive for cross-lingual transfer performance. For instance, syntactic similarity is most predictive for POS tagging and parsing, whereas trigram overlap is the most important predictor for $n$-gram-based topic classification (§5.1);
@@
    Practical implication within studied tasks: Choosing a source language based on a pertinent similarity measure leads to adequate transfer results (§5.2.1);
@@
    Practical implication across studied tasks: When no information about pertinent similarity measures is available for a given task, choosing a source language based on findings for a conceptually similar experiment is a relatively safe choice (§5.2.2).
@@ 2 Related Work
Table 1 provides an overview of related work, highlighting an apparent trade-off between including many languages or including multiple tasks. An important difference that distinguishes our work from prior work is the focus on both. We argue that including a large number of both source and target languages is important for ensuring a relatively balanced distribution of languages. [^1] I.e., such that there is not mainly one cluster of (European) languages that are similar to each other, and most other languages are both dissimilar to that cluster and to each other. In comparison to the studies that come closest to our work in terms of number of included source and target languages de Vries et al. (2022); Samardžić et al. (2022), we include a larger variety of tasks, allowing us to draw conclusions across task boundaries.
Philippy et al. (2023) survey contributing factors for cross-lingual transfer, including most of the works in Table 1. An important take-away from their work is that prior research presents contradicting findings. E.g., some studies find lexical overlap to correlate more strongly with token rather than sentence classification tasks Srinivasan et al. (2021), whereas others find the opposite Ahuja et al. (2022). Similar contradictory examples are presented for different linguistic similarity metrics, pre-training dataset size, and model architecture.
Based on this insight, Philippy et al. make recommendations for follow-up work: (i) focus on real natural languages (instead of synthetic ones), (ii) examine the interaction between different contributing factors, (iii) focus on many languages, (iv) focus on linguistic features when selecting training languages, and (v) focus on generative tasks, given the success of generative models. We focus on the first four recommendations in this work. We include three classification tasks (POS tagging, dependency parsing, and topic classification), motivated by the availability of train and test data in a large number of languages for these tasks.
[…]
Table 1: Related work focusing on zero-shot transfer between many languages or on many tasks. Tasks: D=dependency parsing, E=entity linking, I=natural language inference, M=machine translation, N=named entity recognition, P=part-of-speech tagging, Q=question answering, S=sentence retrieval, T=topic classification.
@@ 3 Methodology
We run transfer experiments on two word-level grammatical tasks with word-level annotations (§3.1) and a sentence-level topic classification task (§3.2). Of the 266 languages in our experiments, 55 training and 84 test languages are shared between all tasks. Appendix §A contains details on all languages and datasets in this study. §3.3 introduces our similarity measures.
@@ 3.1 Grammatical tasks
We use the UDPipe 2 models (Straka, 2018, 2023) that were trained on UD 2.12 data. Each model is trained on a single treebank. The models have so far only been systematically evaluated on their within-treebank performance, but not cross-lingually. UDPipe 2 combines monolingual character and word embeddings [^2] The embeddings are trained on the available training treebanks for a given language. with multilingual embeddings derived from mBERT Devlin et al. (2019). The models are trained jointly for POS tagging, dependency parsing, lemmatization and morphological feature prediction. We only evaluate on the first two, as many treebanks do not have gold standard labels for the others. UDPipe 2 post-processes the predicted dependencies to ensure that each sentence includes a root dependency that all other nodes (in)directly depend on.
We evaluate POS tagging using accuracy, and dependency parsing using the labeled (LAS) and unlabeled attachment scores (UAS). For LAS, we follow UD’s evaluation scripts and ignore dependency label subtypes.
@@ 3.3 Similarity measures
We include a range of dataset-dependent and -independent similarity measures, which are similar to measures used in related work (§2).
@@
Grambank Skirgård et al. (2023) encodes grammatical information for several thousand languages. Its 195 grammatical features are chosen to allow almost no logical dependencies between the values of different features, i.e. the value of one feature does not logically entail the value of another.
@@
We use lang2vec’s location information (each language is represented as a vector of distances to a number of points on the Earth’s surface) and calculate the Euclidean distances between language vectors. We define geographic proximity (geo) as 1 minus the distance. [^5] We do not use lang2vec’s precomputed distances, as the package erroneously returns the maximum distance for some languages, regardless of the comparison.
@@
We count the number of sentences in each training dataset (size). For the topic classification task, this is trivial as each language has the same number of training samples.
@@
4 Supporting Results
Here we present the results on which the main analyses in §5 build. In §4.1 we present the POS tagging, parsing and topic classification scores within and across languages. In §4.2 we compare how (dis)similar transfer patterns are across tasks and experiments, which helps contextualize the importance of different similarity measures for different experiments and tasks in §5.1.
@@ 4 Supporting Results
We focus on the language level for most of our analyses. For UD, which has nearly twice as many datasets as languages, we find that trends are very stable across datasets of the same language (§E.2).
4.1 Within-language vs. cross-lingual performance per task
@@
As expected, the within-language performances are much higher than the cross-lingual performances. Table 2 shows the mean scores within and across languages. Heatmaps with individual scores can be found in §E.1. Table 9 in §E additionally provides dataset-level scores and comparisons to scores reported in related work. Most notably, we find that our POS tagging results are similar but not identical to the ones that de Vries et al. (2022) obtained on UD v2.8 with fine-tuned XLM-R models for 65 training and 105 test languages. When we select the v2.8 languages and use language-level averages, our results are similar (Pearson’s $r$ and Spearman’s $\rho=0.73$, $p<0.0001$) – while the model choice (incl. input representations) matters for the resulting patterns, many transfer trends are similar.
Within-lang. Cross-ling. POS (accuracy) 93.4 6.8 43.0 20.2 Parsing (UAS) 84.8 10.1 36.8 18.8 Parsing (LAS) 78.4 13.7 20.6 17.0 topics-base (acc.) 66.7 14.1 20.1 8.8 topics-translit (acc.) 66.7 10.8 22.5 8.2 topics-mbert (acc.) 58.2 19.9 42.6 20.4
Table 2: Mean scores (in %) of evaluations within vs. across languages (standard deviations in subscripts).
@@
For topic classification, we compare the different input representations. The within-language performance is higher for the monolingual $n$-gram-based models than the models with multilingual input representations from mBERT. We hypothesize that this due to high [UNK] token rates for some languages that were not in mBERT’s pre-training data. [^6] Bagheri Nezhad and Agrawal (2024) similarly find that the performance of multilingual models that are monolingually trained and evaluated on SIB-200 relates to how well a language is represented in a model’s pre-training data.
Although topics-base and topics-translit achieve the same within-language accuracy, topics-translit performs slightly better cross-lingually, having the advantage of higher $n$-gram overlap between datasets. [^7] We further analyze the effect of writing systems in §E.3. In our correlation analysis (§5.1), (lack of) character overlap serves as a proxy for writing system differences. In cross-lingual evaluations, topics-mbert benefits from its multilingual input embeddings: its accuracy is about twice as high (42.6%) as for the monolingual models. Generally, transfer works well if both training and test languages are in mBERT’s pre-training data (or closely related to a language that is): The average accuracy is 68.3% if both languages were included in mBERT’s pre-training data, 46.5% if only the target language was included, 36.1% if only the source language was included, and 33.2% if neither was included. This is consistent with results by Adelani et al. (2024), who fine-tuned four XLM-Rlarge models on one high-resource language each and found their transfer results to be very similar to each other.
@@
Figure 2: Different experiments produce different transfer patterns. NLP transfer results for all combinations of training (columns) and test languages (rows). The darker a cell, the higher the score. The three heatmaps for the grammatical tasks are sorted in the same order, and the three heatmaps with the topic classification results are sorted in the same order. The darker diagonal shows the within-language scores. Large, labelled heatmaps are in Appendix §E.
4.2 Comparing transfer patterns across tasks
@@
In Figure 2, we plot heatmaps of the transfer results for the different experiments. Darker colors correspond to higher scores. To save space, we omit the language labels (which can be found in §E.1). We compute correlations between the results, which we summarize below (details in §F.1).
@@
The results of the grammatical tasks are highly correlated with each other, the correlations between other task results are weaker. [^8] Because of the very high correlation between the parsing measures LAS and UAS (Pearson’s $r=0.95$), we focus on LAS in the remainder of the paper as the more extensive of the two: LAS not only measures whether the dependency arcs are correct, but also whether they were labelled correctly. Even when using exactly the same underlying text data and models, evaluating on a related task results in transfer trends that are similar, but not entirely the same: The parsing (LAS) and POS tagging results by identical UDPipe models (except for the classification heads) are correlated with $r=0.86$.
@@
The preprocessing of the input data (e.g., transliteration) and especially the choice of input representation (mono- vs. multilingual) affects the transfer trends: The results of the two $n$-gram-based topic classifiers are highly correlated with each other ($r=0.68$), but not with the results of the mBERT-based set-up ($r=0.28$ and $0.36$). Set-ups that involve multilingual representations are more highly correlated. For the languages that appear both in UD and SIB-200, the results of the grammatical tasks are most strongly correlated with those of the mBERT-based topic classification models ($r=0.64$ for POS tagging, $r=0.58$ for LAS).
5 Main Results and Analysis
Here, we investigate which factors correlate with overall transfer performance (§5.1). Then, we explore what this means for selecting a source language for cross-lingual transfer (§5.2).
5.1 How do the similarity measures correlate with the tasks and input representations?
@@
We calculate the correlations between the similarity measures and the NLP task results. To account for differences between test sets, we calculate correlations (Pearson’s r) between similarity measures and transfer results on a test language level. We compare the performance of different NLP systems on the same test language, but we do not compare the performance of a single system across multiple test sets. This decision is motivated especially by the challenges of comparing parsing performance across treebanks. Sentence length influences parsing difficulty as it determines the available search space (cf. McDonald and Nivre, 2011, Choi et al., 2015), and morphological differences between languages hamper the comparability of parsers Nivre and Fang (2017).
@@
We analyze overall correlation scores by averaging across test languages. [^9] We treat correlation coefficients with p-values of at least 0.05 as 0, and exclude items where we could not calculate similarity scores due to missing linguistic information. We use language-level averages so that those languages with multiple training and test datasets do not have artificially high correlations, and languages with multiple test datasets do not have greater influence on the overall averages. However, if we instead consider treebank-level correlations, we observe that test sets that belong to the same language show very similar correlation patterns and scores (§E.2).
@@
The correlations between task results and similarity measures vary across our experiments. Figure 3 shows the correlations, which we summarize below. Unaggregated correlations and 95% confidence intervals of the mean scores can be found in §F.3. These unaggregated correlations show that, while there are some outliers, the overall trends we report below accurately reflect the language-level trends.
@@
Figure 3: Mean correlation scores between task results and similarity measures. “Word*” = overlap between words (wor; UD tasks) trigrams (tri; topics-base/translit), and subword tokens (swt; topics-mbert). Dotted lines are added for intelligibility.
@@
We observe some common trends across experiments: Training dataset size does not matter much: it is the same across all topic classification experiments, and the correlations between training set size and performance in the grammatical tasks are close to zero. The phonological and phonetic similarity measures (pho, inv) also generally have low correlation scores. [^10] Philippy et al. (2023) find that phonological similarity is more important for token- than sentence-level tasks. This is different in our study, where the correlations with inv are higher for the $n$-gram-based topic classification models. We assume that this is due to phonological similarity being moderately correlated with string similarity measures like trigram overlap (§C.2), which is in turn strongly correlated with the topic classification results.
@@
The strongest predictor for parsing performance is syntactic similarity (syn) as determined by lang2vec ($r$avg_syn$=0.57$), followed by the similarity of Grambank features (gb; $r$avg_gb$=0.42$). These correlations are even stronger when we only consider the test languages that mBERT was pretrained on ($r$avg_syn$=0.69$ and $r$avg_gb$=0.60$, respectively; Appendix §F.3).
The POS tagging outputs show correlation patterns similar to the ones for parsing, albeit weaker. Although the correlation strengths are generally weaker for POS tagging, e.g., the strongest predictor is lang2vec’s syntactic similarity (syn) with $r$avg_syn$=0.37$, word overlap (wor) is as good a predictor for POS accuracy as for LAS ($r$avg_wor=0.33 and $r$avg_wor=0.34). Again, the correlation strengths are higher when only considering test languages that were in mBERT’s pre-training data ($r$avg_syn$=0.46$). [^11] The correlation strengths are also higher when we analyze de Vries et al.’s results ($r$avg_syn$=0.46$). We hypothesize that this is due to the smaller amount of test languages in their experiments, a higher proportion of which is in XLM-R’s pre-training data. The correlations we observe for the grammatical task’s transfer results partially align with prior research: Lauscher et al. (2020), Samardžić et al. (2022) and Pires et al. (2019) also find syntactic similarity to be important, and Lin et al. (2019) and Xia et al. (2020) also find word overlap to be a good predictor for POS tagging. However, contrary to our results, the latter two find syntactic similarity to be relatively unimportant for both tasks. We hypothesize that such inconsistencies are due to differences in the language sets studied and other experimental differences (e.g., Lin et al. (2019) add data from the test language to the training set when possible).
@@
The results of the $n$-gram-based models are most highly correlated with measures of string similarity and lexical similarity. This is expected based on their input representations. For topics-base, the trigram overlap (tri) between training and test data shows the highest correlation ($r$avg_tri=0.65), followed by character overlap (chr; $r$avg_chr=0.49) and lexical similarity (lex; $r$avg_lex=0.49). Trigram overlap and lexical similarity (and the highly correlated genetic relatedness, gen) are also the strongest predictors for topics-translit ($r$avg_tri=0.55, $r$avg_lex=0.61, $r$avg_gen=0.55). However, for this model, character overlap only plays a very small role ($r$avg_chr=0.11), as all datasets use the same character inventory due to the transliteration.
@@
For the model using mBERT representations (topics-mbert), none of the correlations are strong. The correlations peak at $r$avg_chr=0.27 for character overlap (chr). Model performance depends instead on the inclusion of the source and especially target languages in mBERT’s pre-training data (§4.1).
@@
The transfer results cannot be predicted by any one factor alone. We fit a linear mixed effects model for each experiment, with the NLP score as the dependent variable and the source and target languages as random effects. The fixed effects are the similarity scores, whether the training and test data use the same writing systems, and whether the test language was included in mBERT’s pre-training data. We use R R Core Team (2024) and the lme4 package Bates et al. (2015). This analysis involves a reduced set of languages, as for many language pairs in our data at least one similarity metric is not defined and we do not perform any data imputation. The analysis shows similar trends to the correlations described above. Additionally, for each experiment, most of the variables in our analyses are significant predictors of the transfer score (details in §F.2). This even sometimes applies to measures that capture similarity at the same linguistic level: e.g., both grammatical similarity measures (syn, gb) are independently among the highest predictors for parsing performance. The finding that multiple measures contribute to predicting the transfer score confirms previous work Lin et al. (2019); de Vries et al. (2022).
5.2 Practical implications
So far we have investigated correlations on a global level. In this section we investigate how these findings can be used in practice: Do the overall correlation patterns also apply when picking a source language for a given target language according to simple heuristics derived from the global patterns?
We intentionally pick very simple heuristics, as we believe them to be more realistic to how practitioners choose source language candidates in practice. Additionally, we choose them to be easily generalizable and not be constrained to the languages in our experiments. [^12] For instance, while LangRank Lin et al. (2019) is an easy-to-use tool for transfer recommendations, it is limited to the tasks and languages it was trained on and thus recommends one of 30 possible source languages for dependency parsing or one of 60 for POS tagging.
5.2.1 Should I pick the most similar language according to one similarity measure?